Operations Report – January 17th

Crew 203 Operations Report 17-JAN-2019

SOL: 5

Name of person filing report: Freddy Castaneda

Non-nominal systems: Suit #8 not charging, pending to review. Generator

Notes on non-nominal systems: Generator had to be jump-started, then later began running on low power. Troubleshooting in progress

Generator (hours run): 17hr 45min; Turned on last night (16Jan2019) at 16:30; Turned off this morning (17Jan2019) at 10:15; Turned on tonight (17Jan2019) at 18:15

Solar SOC – Turned on (16Jan2019) 95%; Turned off (16Jan2019) 100%; Turned on (17Jan2019) 74%

Diesel Reading – Not read, EVA in progress.

Propane Reading – Not read, EVA in progress.

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – Over 70%; About 375 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 76%; 227 gallons

Water (loft) – At level marker 7

Static to Loft Pump used – Yes; At 9:20 to refill tank

Water Meter: 01401702

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes, used on a EVA to Kissing Camel

Hours: 68.2

Beginning charge: 100%

Ending charge: 68%

Currently charging: Yes

Opportunity rover used: Not used

Hours: 45.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes, used on a EVA to Kissing Camel

Hours: 70.0

Beginning charge: 100%

Ending charge: 69%

Currently charging: Yes

Notes on rovers: Spirit was flashing a red maintenance light during EVA, now is charging and pending to review.

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: All ATV´s were gas filled, we have 10 gallons of ATV gas.

HabCar used and why, where? Not used. We were informed that the Hab Car washer fluid container has a leak, it might not be used, a spray bottle with washer fluid for use will be created by Shannon and Atila to cover the need.

CrewCar used and why, where? Off campus

General notes and comments: The generator had to be jump started tonight. No idea why.

Summary of internet: Nothing to report

Summary of suits and radios: Suit #3 still not charge, now Suit #8 does not charge either, pending for review.

Summary of Hab operations: The commanders room door lock was fixed, now he can hide from us if he wants.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing to report